12 mole of helium is contained in a container at S.T.P. How much heatenergy is needed to double the pressure of the gas, keeping the volumeconstant? Heat capacity of gas is 3 J g–1 K–1.

Grade:11

1 Answers

Arun
6 years ago
n = ½
Cv = 3J/gm-K
M = 4
Cv = MCv = 12 J/gm-K      M = molecular mass
P2/P1 = T2/T1 = 2
Delta T = T2- T1 = 273 K
Delta Q = n Cv (Delta T)
 On solving
Delta Q = 1638 J
 
Regards
Arun (askIITians forum expert)